In a triangle, the measure of he middle angle is 17 degrees more than twice the measure of the smallest angle. The largest angle is 55 degrees more than the measure of the smallest angle. What are the measures of the three angles?
Found 3 solutions by Paul, troyapplehelen, checkley71:
Answer by Paul(988)

Triangle adds up to 180 degrees:
Let one side be x
Let middle side be 2x+17
Let the largest side be x+55
Equation:
(x)+(2x+17)+(x+55)=180
4x+72=180
4x=108
x=27

2(27)+17=71
27+55=82.
Hence, the three sides are 27, 71 and 82 degrees.
Paul.

Answer by troyapplehelen(46)

1. the smallest is x
2. then the mid is 2x+17
3. the largest is x+55
4. the sum of the interior angle in a triangle is 180
so, x+(2x+17)+(x+55)= 180
4x+72=180
4x=180, x=27 then 2x+17(the mid) = 81 and the largest is 82.

Answer by checkley71(8403)

(17+2X)+(X+55)+X=180 OR 17+2X+X+55+X=180 OR 4X+72=180 OR 4X=108 OR X=27
PROOF (17+54)+(82)+27=180 OR 180=180